OpenStudy (anonymous):

Solve: 3(x + 4) − 4(x + 3) = 7x + 16

OpenStudy (unklerhaukus):

expand the brackets

OpenStudy (anonymous):

do it like this 3(x+4)-4(x+3) = 7x+16 = 3x+12 -4x -12=7x +16 get a like terms together like this 3x-4x-7x=16 the 12-12 is zero so no need. -8x=16 get x my itself divide by -8/16 so you have X= -2 hopes this helps